OpenStudy (anonymous):

need help on solving 3/y+7/2y=13

OpenStudy (anonymous):

\[\frac{3}{y}+\frac{7}{2y}=13\] multiply both sides by \(2y\) and start with \(6+7=23y\)

OpenStudy (anonymous):

I thought u had to multiply by 2 for 3/y to get the denominator the same

OpenStudy (anonymous):

yeah that is why i got \(6\) for the first number, not \(3\)

OpenStudy (anonymous):

I did that but what do you do afterwards

OpenStudy (anonymous):

also i made a typo it should be \[6+7=26y\]

OpenStudy (anonymous):

what do you do now? what is \(6+7\) ?

OpenStudy (anonymous):

its 13 but what about the 2y

OpenStudy (anonymous):

i see i have confused you lets start at the beginning the common denominator for \(\frac{3}{y}\) and \(\frac{7}{2y}\) is \(2y\)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

start with your equation \[\frac{3}{y}+\frac{7}{2y}=13\] and multiply both sides by \(2y\) cancelling as you go

OpenStudy (anonymous):

\[2y(\frac{3}{y}+\frac{7}{2y})=13\times 2y\]

OpenStudy (anonymous):

\[2y(\frac{3}{y}+\frac{7}{2y})=13\times 2y\]\[2y\times \frac{3}{y}+2y\times \frac{7}{2y})=13\times 2y\] \[6+7=26y\]

OpenStudy (anonymous):

the 6/2y+7/2y=13 from there I have to multiply by 2y on each side to get rid of it

OpenStudy (anonymous):

\[2\cancel y\times \frac{3}{\cancel y}+\cancel{2y}\times \frac{7}{\cancel{2y}}=13\times 2y\\ 6+7=26y\]

OpenStudy (anonymous):

ok I understand

OpenStudy (anonymous):

good!

OpenStudy (domebotnos):

Can u look at my question please? http://openstudy.com/study#/updates/53a4e625e4b0ffdda15dfaef